OpenStudy (anonymous):

Solid xenon hexafluoride is prepared by allowing fluorine gas and xenon gas to react. Write a balanced equation for this reaction. Using this equation, tell me how many grams of xenon are required to form 20.0 g xenon hexafluoride.

OpenStudy (gebooors):

Xe (g) + 3 F2 (g) -> XeF6 (s) Xe does not exist in molecular form. Calculate molar mass M of XeF6. Then calculate amount of substance n =m/ M M = 20 g. Then you see n(Xe) = n(XeF6). Mass of Xenon m = n x M(Xe)